OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x3 + 8x Odd Even Neither

OpenStudy (anonymous):

odd

OpenStudy (anonymous):

i have a couple other questions if you don't mind helping me ? (:

OpenStudy (anonymous):

ill see what I can do

OpenStudy (anonymous):

Perform the requested operation or operations. Find the domain of each. f(x) = 7x + 6, g(x) = 4x2 Find (f + g)(x). 7x + 6 + 4x2 28x3 + 24x 7x + 6 - 4x2 7x+6/4x^2 @Jpenrod

OpenStudy (anonymous):

this might be out of my reach of intelligences

OpenStudy (anonymous):

i've been thinking the same thing since i started doing this ...

OpenStudy (anonymous):

do you know how to find vertical and horizontal asymptotes ?

OpenStudy (anonymous):

give me some of the asymptotes it might remind me of what they are

OpenStudy (anonymous):

For the given function, find the vertical and horizontal asymptote(s) (if there are any). f(x) = x^2+3/x^2-9 a) x = 3, x = -3, y = 1 b) x = 9, y = 3 c) None d) x = 9, y = 1

OpenStudy (anonymous):

http://www.mathway.com/problem/ODUyNDY1NjE tell me if this helps